if 6 people do a certain job in 8 hours how long will it take 16 people working at the same speed?
Answer by stanbon(75887)

if 6 people do a certain job in 8 hours how long will it take 16 people working at the same speed?
---
# of hrs and # of people are indirectly related.
----
h = k/p
---
Solve for "k" using "if 6 people do a certain job in 8 hours".
8 = k/6
k = 48
----
Equation:
h = 48/p
---
how long will it take 16 people working at the same speed?
h = 48/16
h = 3 hrs.
===============
Cheers,
Stan H.
===============